OpenStudy (anonymous):

Write the equation 5x + 4y = 10 so that y is a function of x. You must show work in order to receive full credit. A. y= 5/4x + 2.5 B. y= 1.2x - 5/2 C. y= -5/4 - 5/2 D. y= -5/4x + 5/2

OpenStudy (anonymous):

The question is essentially asking you to convert the equation, which is in standard (ax + by = c) form, into slope intercept form (y = mx + b; m being the slope and b being the y intercept). The slope of an equation in standard form is -a/b. In this case it would be -5/4. The y intercept of an equation in standard form is c/b. In this case it would be 10 / 5 = 2. The other way to do it (i.e. how that shortcut was derived) is through algebra. Isolate y on the left side of the equation by subtracting 5x from both sides and then dividing both sides by four.

OpenStudy (anonymous):

So does (x) equal 2 ?

OpenStudy (anonymous):

The question isn't asking you to solve for x. However, if you were to isolate x on the left side, x = 2 - (4/5)y is the answer, not x = 2.